# Critical Evaluation of Studies Alleging Evidence for Technosignatures in the POSS1-E Photographic Plates

### v1

@@ I Introduction
Villarroel et al. (2021, 2022) described searches in astronomical data records from the pre-Sputnik era for local technosignatures in the form of glinting artificial satellites. In the present study, we perform a critical examination of the data, analyses, and results presented in recent papers by Villarroel et al. (2025a) and Bruehl and Villarroel (2025) that build on this previous work. Both papers use datasets of nominally unidentified features that derive from datasets generated by Solano et al. (2022) using two sets of complete digital scans of the first-epoch Palomar Observatory Sky Survey (POSS1, 1949-1957): (i) the Space Telescope Science Institute’s Digital Sky Survey [^1] https://archive.stsci.edu/dss/acknowledging.html (DSS) and (ii) the SuperCOSMOS survey (Hambly et al., 2001). The first paper (Villarroel et al., 2025a) suggests that some fraction of the features may represent objects reflecting sunlight in near-Earth space. This conclusion is heavily based on an alleged deficit of features in the Earth’s shadow. Previous work was also concerned with the search for statistically significant alignments and clusters of the purportedly unidentified POSS1 features (Solano et al., 2024; Villarroel et al., 2021, 2022). The second study (Bruehl and Villarroel, 2025) reports a correlation between the days when features were recorded and the timing of nuclear weapons tests as well as when UAP sightings were reported.
@@
Regarding the SPFs used in Villarroel et al. (2025a) and Bruehl and Villarroel (2025) to estimate correlations with the terrestrial shadow and nuclear tests, our work suggests that (i) a significant fraction (at least 91%) of these were part of a set removed by Solano et al. (2022) for residing within 5′′ of catalog stars; (ii) a potentially significant but undetermined amount are at risk of being star-like plate emulsion artifacts; and (iii) up to 40% are at risk of being scan artifacts of the type removed by Solano et al. (2022) to obtain their most vetted dataset. Villarroel et al. (2025a) acknowledged222e.g., See Section 8, pg 16 of Villarroel et al. (2025a): “this sample… is expected to contain a substantial number of false positives, including clustered artifacts such as edge fingerprints or other plate defects that contaminate our sample.” the possibility that a significant fraction of SPFs in the analyzed datasets are the result of emulsion defects, degradation, and contamination, and did not formally validate these datasets in order to rule this out.
In Section II, we briefly review the literature about a two-decade search in archival plates for then-novel optical transients related to gamma ray bursts (GRBs), which revealed common star-like emulsion defects and the difficulties in distinguishing these from optical flashes. In Section III, we define and describe the datasets under consideration while pointing out significant ambiguities and inconsistencies, as well as caveats associated with their principal uses in Villarroel et al. (2025a) and Bruehl and Villarroel (2025). Section IV addresses the spatial distribution of SPFs between plates, within plates, and across the sky, suggesting plate artifacts as their likely origin. We also examine the SPFs reported to occur within sporadic linear clusters in Villarroel et al. (2025a), finding that a third of these were discarded by Solano et al. (2022). Section V critically examines the purported deficit of SPFs in Earth’s shadow reported in Villarroel et al. (2025a) and offered as evidence that these represent glinting artificial satellites in the pre-Sputnik era. Using the most vetted dataset in Solano et al. (2022), we find no evidence of a deficit. Section VI addresses the alleged correlation between SPFs and the timing of nuclear tests as reported in Bruehl and Villarroel (2025), finding that this correlation is almost entirely determined by the Palomar telescope observation schedule. The discussion in Section VII focuses on major challenges to the analyses and conclusions in both of the studies we have examined, including the description a circular argument used to justify confidence in the conclusions as well as the validity of measurements.
@@ II Historical searches for optical transients in photographic plates
Villarroel et al. (2025a) and Bruehl and Villarroel (2025) analyzed SPF datasets that we describe in Section III and were motivated by questions developed in Solano et al. (2022), Solano et al. (2024), Villarroel et al. (2021) and Villarroel et al. (2022). The goal of their work is to uncover evidence of technosignatures in the form of unidentified optical transients in pre-Sputnik photographic plates acquired by the National Geographic POSS1 survey (Minkowski and Abell, 1963). The validity of this entire body of work rests on the premise that the selected features are in fact images of real optical transients distinct from stars and other celestial bodies, and are not plate artifacts or other defects of spurious origin unrelated to objects in the sky.
The confident identification of optical transients (short-lived optical emissions) in archival photographic plates was the goal of a two-decade-long effort to identify the first optical flash counterpart of a gamma-ray burst (GRB). GRB optical counterparts were postulated as a new and unproven source of optical transients (OTs), similar to the role played by non-anthropogenic technosignatures in the work of Villarroel and colleagues, except that the GRBs were an independently verified phenomenon with estimates of position in the sky. Work began in 1974 (Grindlay et al., 1974), but the first gamma-ray OT candidate was announced in 1981 after searching 4,135 physical plates from Harvard University Observatory spanning 1889-1953 and 1970-1979 (Schaefer, 1981). Researchers then combed the physical plate collections of the Harvard, Sonneberg, Ondfejov, and Bamberg observatories ($\sim$910,000 plates; Hudec et al. (1994)) and attempted to determine if any of the transient, star-like OT candidates found in the photographic emulsion of archival plates were: (i) the result of an optical flash; (ii) not from any known atmospheric or celestial event; and (iii) a counterpart to a GRB at a previously known location (see for example: Greiner et al. (1987); Zytkow (1990); Schaefer (1990); Greiner (1992); Schaefer (1994); Greiner and Moskalenko (1994); Hudec (1993); Vrba et al. (1995); Fishman and Meegan (1995)).
@@
Initial assessments about the validity of each star-like OT candidate centered on techniques that helped distinguish between a star-like plate artifact and a true optical transient. Ideally, a blink test was performed (which finds differences between two or more simultaneous photographs containing the same candidate). Without this, it was necessary to conduct an analysis of the physical plate itself using microscopy (500x) and microdensitometry (to obtain photodensity image profiles) with transmitted and reflected light (to look at grain structure in the emulsion), as well as CCD scanning techniques. These techniques were used in an effort to find: (i) a valid star-like 3-D structure in developed silver grains within the plate emulsion, and evidence of short duration (evidence of an optical flash); (ii) coma distortion (evidence of optical path); (iii) halos from atmospheric fog (evidence of optical path) (Hudec et al., 1994); and (iv) lack of a trailing (caused by telescope tracking errors) might indicate a short duration optical flash, but could also be a plate defect (Schaefer, 1990). In the absence of multiple simultaneous plates, OT candidates with the supplementary support provided by these techniques were more likely to be accepted by scientific peers (Hudec, 1993). However, several groups applied these techniques to the same plate and arrived at different conclusions (Hudec, 1993; Vrba, 1996). By 1995, several OT candidates were still being championed as real optical flashes, although not unequivocally associated with GRBs (Vrba, 1996).
@@
Digitization of the DSS POSS1 plates began in 1992 and they became available on CD-ROM in 1995; the SuperCosmos digitization of POSS1 became available in 2000. The physical plate set of SuperCosmos copy negatives is available for examination at the Royal Observatory, Edinburgh (Hambly and Blair, 2024); many surveys have slightly inferior originals available to examine (Morgan, 1995).
The DSS and SuperCosmos digital scans used in Solano et al. (2022) and Villarroel et al. (2025a) added additional pathways for processing-related artifacts to enter the picture, such as images of dust particles and emulsion backing fragments.
In light of the challenging results from two decades of GRB OT studies, and the very few optical transient candidates found, it is not clear how a search for alignments and correlations of optical flashes from non-anthropogenic technosignatures proceeds without first independently verifying that the detections in the dataset(s) represent optical flashes.
Expressing similar concerns, Hambly and Blair (2024) examined the intensity profile statistics (position, brightness, and morphology) of all objects on the POSS1-E plate E0070 containing the nine transients allegedly identified by Villarroel et al. (2021). They found a class of objects they designated as “spurious,” that are statistically distinctive, and morphologically distinguishable from stars and other astronomical objects; e.g., they exhibit a smaller full-width at half maximum (FWHM) on average. They then trained a machine learning model to distinguish between stars, galaxies, and likely spurious (not within $5^{\prime\prime}$ of a Gaia or Pan-STARRS DR1 catalog object) and applied it to plate E0070, finding $\sim 45,000$ stars, $\sim 27,000$ galaxies, and $\sim 8000$ spurious objects (including the nine “transients” from Villarroel et al. (2025a)). Their so-called spurious objects nominally belong to the same class as the so-called transient dataset (N=298,165) defined by Solano et al. (2022) and used in Villarroel et al. (2025a). Villarroel et al. (2025b) have since acknowledged that many of the unidentified SPFs exhibit a smaller FWHM when compared with stars, but have countered that this is not unexpected for short-lived flashes from objects in high-altitude orbits, during a long-duration (50 min) exposure. The following questions remain unanswered after Hambly and Blair (2024): (i) whether optical flashes exhibit smaller FWHM on POSS1 plates; (ii) whether any of the 8,000 features selected on plate E0070 by Hambly and Blair (2024) represent an optical flash; and (iii) whether some categories of plate artifacts exhibit a smaller FWHM when compared to stars, on average.
@@ III Data sources
In this section, we describe (i) the publicly available datasets used for the analyses in this paper; and (ii) the publicly known properties of the related and as-yet unpublished datasets used by Bruehl and Villarroel (2025) and Villarroel et al. (2025a). Access to the unpublished datasets is not necessary for our evaluation of the analyses and claims in the papers under scrutiny. The relationships between these published and unpublished datasets are illustrated in Figure 1 and defined in Table 1, where “SVR22” stands for Solano, Villarroel, and Rodrigo (2022), “VEA25” stands for Villarroel et al. (2025a), and “BV25” stands for Bruehl and Villarroel (2025). We use our own naming convention when referring to datasets (see the symbol (“Sym.”) column in Table 1), which are not used in the papers we have cited.
@@ III.1 Definitions
The POSS1-E plates comprise 937 individual exposures using the red-sensitive Kodak 103a-E plate emulsion, whose sensitivity peaks at 640 nm with a limiting magnitude of 20 (Minkowski and Abell, 1963). This optical survey ran from November 11, 1949 to December 10, 1958 and covered the whole sky north of declination -33. This survey captured exposures for 292 plates with footprints residing fully in the southern celestial hemisphere; 584 plates with footprints residing fully in the northern celestial hemisphere; and 61 with footprints that overlap the celestial equator. The DSS and SuperCOSMOS scans were made from two independent sets of glass copy negative plates made from a single set of glass copy positive plates, in turn made from the original glass negatives (Hambly and Blair, 2024).
@@
To efficiently remove known stars while accounting for proper motion between surveys, SPFs in $A$ within 5′′ of catalog objects from Gaia DR3 (Vallenari et al., 2023) and Pan-STARRS DR2 (Chambers et al., 2019) were removed from $A$ (along with an unknown number of proximate SPFs that were not stars) to obtain the subset $S$ (N = 298,165). With the exception of one plate, plates whose footprints reside entirely within the southern celestial hemisphere were never sampled, or else SPFs from those plates were removed at this stage.
@@
For unknown reasons, Villarroel et al. (2025a) and Bruehl and Villarroel (2025) did not use $R$. Instead, Villarroel et al. (2025a) provided ambiguous definitions of the four datasets used in its analyses ($S$, $V$, $V^{\prime}$, and $V^{\prime\prime}$), which are generally not concordant with definitions in Solano et al. (2022). Set $S$ is tied to its definition in Solano et al. (2022) by the reported size ($N=298,165$).
@@
Although Bruehl and Villarroel (2025) cited Solano et al. (2022) as the source of its primary dataset, $R$ was not used in this case, either. This was instead the set $V$ from Villarroel et al. (2025a), based on the reported size ($N=107,875$).
Our study makes use of the following SPF datasets, all of which are publicly available: (i) set $W$ ($N=171,753$), consisting of SPFs near NeoWISE object locations; (ii) set $R$, the most highly vetted subset from Solano et al. (2022); and (iii) set $M$, which contains the nearly 90 million SPFs recorded in the POSS1-E and POSS1-O plates by the MAPS project (Pennington et al., 1993; Cabanela et al., 2003). MAPS discarded all features that did not appear in both the E (red band) and O (blue band) plates, in order to avoid recording scratches and other plate artifacts; the Galactic plane was also occluded ($|b|>20^{\circ}$) to avoid features that blend together in crowded fields. We chose the MAPS project digitization for comparison because it was independent of the DSS and SuperCOSMOS digitizations and because it records only very confident detections of celestial objects. We have also used the POSS1-E meta data (STScI, 2026), as well as POSS1-E images in DSS and SuperCosmos scans (shown in Figures 2 and 9) and Pan-STARRS images (shown in Figure 2). Finally, we use the three sources cited in Bruehl and Villarroel (2025) for nuclear test dates (NNSS, 2023; Wikipedia, 2025; CHERCA, 2025).
Figure 1 illustrates the relationship between the $A$, $S$, $W$, $P$, $V$, and $R$ datasets defined in Table 1, which are defined or implied in Solano et al. (2022), Villarroel et al. (2025a), and Bruehl and Villarroel (2025). This also illustrates the ambiguity (dashed line and “?”) concerning the amount of overlap between the publicly available set $R$, which is the most vetted Solano et al. (2022) dataset, and $V$, which has not been released as of the time of this writing. We stress that the relationship between $R$ and $V$ has not been disclosed, and that access to $V$ is not necessary for our evaluation of the analysis and claims in Villarroel et al. (2025a) and Bruehl and Villarroel (2025).
III.2 Principal uses and additional properties

### v2

@@ I Introduction
Villarroel et al. (2021, 2022) described searches in astronomical data records from the pre-Sputnik era for local technosignatures in the form of glinting artificial satellites. In the present study, we perform a critical examination of the data, analyses, and results presented in recent papers by Villarroel et al. (2025a) and Bruehl and Villarroel (2025) that build on this previous work. Both papers use datasets of nominally unidentified features that derive from datasets generated by Solano et al. (2022) using two sets of complete digital scans of the first-epoch Palomar Observatory Sky Survey (POSS1, 1949-1957): (i) the Space Telescope Science Institute’s Digital Sky Survey [^1] https://archive.stsci.edu/dss/acknowledging.html (DSS) and (ii) the SuperCOSMOS survey (Hambly et al., 2001). The first paper (Villarroel et al., 2025a) suggests that some fraction of the features may represent objects reflecting sunlight in near-Earth space. This conclusion is heavily based on an alleged deficit of features in the Earth’s shadow. A set of candidate linear clusters is presented in this work; previous work was also concerned with the search for statistically significant alignments and clusters of the purportedly unidentified POSS1 features (Solano et al., 2024; Villarroel et al., 2021, 2022). The second study (Bruehl and Villarroel, 2025) reports a correlation between the days when features were recorded and the timing of nuclear weapons tests as well as when UAP sightings were reported.
@@
Regarding the SPFs used in Villarroel et al. (2025a) and Bruehl and Villarroel (2025) to estimate correlations with the terrestrial shadow and nuclear tests, our work suggests that (i) a significant fraction (at least 91%) of these belong to a set removed by Solano et al. (2022) for residing within 5′′ of catalog stars; (ii) a potentially significant but undetermined amount are at risk of being star-like plate emulsion artifacts; and (iii) up to 40% are at risk of being scan artifacts of the type removed by Solano et al. (2022) to obtain their most vetted dataset. Villarroel et al. (2025a) acknowledged222e.g., See Section 8, pg 16 of Villarroel et al. (2025a): “this sample… is expected to contain a substantial number of false positives, including clustered artifacts such as edge fingerprints or other plate defects that contaminate our sample.” the possibility that a significant fraction of SPFs in the analyzed datasets are the result of emulsion defects, degradation, and contamination, and did not formally validate these datasets in order to rule this out.
In Section II, we briefly review the literature about a two-decade search in archival plates for then-novel optical transients related to gamma ray bursts (GRBs), which revealed common star-like emulsion defects and the difficulties in distinguishing these from optical flashes. In Section III, we define and describe the datasets under consideration while pointing out significant ambiguities and inconsistencies, as well as caveats associated with their principal uses in Villarroel et al. (2025a) and Bruehl and Villarroel (2025). Section IV addresses the spatial distribution of SPFs between plates, within plates, and across the sky, suggesting plate artifacts as their likely origin. We also examine the SPFs reported to occur within sporadic linear clusters in Villarroel et al. (2025a), finding that a third of these SPFs were discarded by Solano et al. (2022). Section V critically examines the purported deficit of SPFs in Earth’s shadow reported in Villarroel et al. (2025a) and offered as evidence that these represent glinting artificial satellites in the pre-Sputnik era. Using the most vetted dataset in Solano et al. (2022), we find no evidence of a deficit. Section VI addresses the alleged correlation between SPFs and the timing of nuclear tests as reported in Bruehl and Villarroel (2025), finding that this correlation is almost entirely determined by the Palomar telescope observation schedule. The discussion in Section VII focuses on major challenges to the analyses and conclusions in both of the studies we have examined, including the description of a circular argument used to justify confidence in the conclusions as well as the validity of measurements.
@@ II Historical searches for optical transients in photographic plates
Villarroel et al. (2025a) and Bruehl and Villarroel (2025) analyzed SPF datasets that we describe in Section III and were motivated by questions developed in Solano et al. (2022), Solano et al. (2024), Villarroel et al. (2021) and Villarroel et al. (2022). The goal of their work is to uncover evidence of technosignatures in the form of unidentified optical transients in pre-Sputnik photographic plates acquired by the National Geographic POSS1 survey (Minkowski and Abell, 1963). The validity of this entire body of work rests on the premise that the selected features are in fact images of real optical transients (short-lived optical emissions) distinct from stars and other celestial bodies, and are not plate artifacts or other defects of spurious origin unrelated to objects in the sky.
Fifty years ago, a similar study sought to confidently identify optical transients in archival photographic plates (Grindlay et al., 1974). The goal of what became a two-decade-long effort was to identify the first optical flash counterpart of a gamma-ray burst (GRB). GRB optical counterparts were postulated as a new and unproven source of optical transients (OTs), similar to the role played by non-anthropogenic technosignatures in the work of Villarroel and colleagues, except that the GRBs were an independently verified phenomenon with estimates of position in the sky. Work began in 1974 (Grindlay et al., 1974), but the first gamma-ray OT candidate was announced in 1981 after searching 4,135 physical plates from Harvard University Observatory spanning 1889-1953 and 1970-1979 (Schaefer, 1981). Researchers then combed the physical plate collections of the Harvard, Sonneberg, Ondfejov, and Bamberg observatories ($\sim$910,000 plates; Hudec et al. (1994)) and attempted to determine if any of the transient, star-like OT candidates found in the photographic emulsion of archival plates were: (i) the result of an optical flash; (ii) not from any known atmospheric or celestial event; and (iii) a counterpart to a GRB at a previously known location (see for example: Greiner et al. (1987); Zytkow (1990); Schaefer (1990); Greiner (1992); Schaefer (1994); Greiner and Moskalenko (1994); Hudec (1993); Vrba et al. (1995); Fishman and Meegan (1995)).
@@
[…]
@@
The POSS1-E physical plate set of SuperCosmos copy negatives is housed at the Royal Observatory, Edinburgh (Hambly and Blair, 2024). For many surveys, observations of star fields are repeated if there are quality concerns (e.g. inferior seeing), a practice that generates slightly inferior original plates that can be made available for detailed examination (Morgan, 1995).
The DSS digitization of POSS1-E and -O plates began in 1992 and they became available on CD-ROM in 1995; the SuperCosmos digitization of POSS1-E plates became available in 2000. These DSS and SuperCosmos digital scans used in Solano et al. (2022) and Villarroel et al. (2025a) added additional pathways for processing-related artifacts to enter the picture, such as images of dust particles and emulsion backing fragments.
The challenging results from two decades of GRB OT studies, and the very few optical transient candidates found, suggest that a search for alignments and correlations of optical flashes from non-anthropogenic technosignatures should include verification, independent of the technosignature hypothesis, that the detections in the dataset(s) represent optical flashes.
Expressing similar concerns, Hambly and Blair (2024) examined the intensity profile statistics (position, brightness, and morphology) of all objects on the POSS1-E plate E0070 containing the nine transients allegedly identified by Villarroel et al. (2021). This study found a class of objects they designated as “spurious,” that are statistically distinctive, and morphologically distinguishable from stars and other astronomical objects; e.g., they exhibit a smaller full-width at half maximum (FWHM) on average. They then trained a machine learning model to distinguish between stars, galaxies, and likely spurious (not within $5^{\prime\prime}$ of a Gaia or Pan-STARRS DR1 catalog object) and applied it to plate E0070, finding $\sim 45,000$ stars, $\sim 27,000$ galaxies, and $\sim 8000$ spurious objects (including the nine “transients” from Villarroel et al. (2025a)). Their so-called spurious objects nominally belong to the same class as the so-called transient dataset (N=298,165) defined by Solano et al. (2022) and used in Villarroel et al. (2025a). Villarroel et al. (2025b) has since acknowledged that many of the unidentified SPFs exhibit a smaller FWHM when compared with stars, but have countered that this is not unexpected for short-lived flashes from objects in high-altitude orbits, during a long-duration (50 min) exposure. The following questions remain unanswered after Hambly and Blair (2024): (i) whether optical flashes exhibit smaller FWHM on POSS1-E plates; (ii) whether any of the 8,000 features selected on plate E0070 by Hambly and Blair (2024) represent an optical flash; and (iii) whether some categories of plate artifacts exhibit a smaller FWHM when compared to stars, on average.
@@ III Data sources
In this section, we describe (i) the publicly available datasets used for the analyses in this paper; and (ii) the publicly known properties of the related and as-yet unpublished datasets used by Bruehl and Villarroel (2025) and Villarroel et al. (2025a). Access to the unpublished datasets is not necessary for our evaluation of the analyses and claims in these papers. The relationships between these published and unpublished datasets are illustrated in Figure 1 and defined in Table 1, where “SVR22” stands for Solano, Villarroel, and Rodrigo (2022), “VEA25” stands for Villarroel et al. (2025a), and “BV25” stands for Bruehl and Villarroel (2025). We use our own naming convention when referring to datasets (see the symbol (“Sym.”) column in Table 1), which are not used in the papers we have cited.
@@ III.1 Definitions
The POSS1-E plates comprise 937 individual exposures using the red-sensitive Kodak 103a-E plate emulsion, whose sensitivity peaks at 640 nm with a limiting magnitude of 20 (Minkowski and Abell, 1963). This optical survey ran from November 11, 1949 to December 10, 1958 and covered the whole sky north of declination -34. This survey captured exposures for 292 plates with footprints residing fully in the southern celestial hemisphere; 584 plates with footprints residing fully in the northern celestial hemisphere; and 61 with footprints that overlap the celestial equator. The DSS and SuperCOSMOS POSS1-E scans were made from two independent sets of glass copy negative plates made from a single set of glass copy positive plates, in turn made from the original glass negatives (Hambly and Blair, 2024).
@@
To efficiently remove known stars while accounting for proper motion between surveys, SPFs in $A$ within 5′′ of catalog objects from Gaia DR3 (Vallenari et al., 2023) and Pan-STARRS DR2 (Chambers et al., 2019) were removed from $A$ (along with an unknown number of proximate SPFs that were not stars) to obtain the subset $S$ (N = 298,165). With the exception of one plate, those plates whose footprints reside entirely within the southern celestial hemisphere were never sampled, or else SPFs from those plates were removed at this stage.
@@
For unknown reasons, Villarroel et al. (2025a) and Bruehl and Villarroel (2025) did not use $R$. Villarroel et al. (2025a) used four datasets in its analyses ($S$, $V$, $V^{\prime}$, and $V^{\prime\prime}$), which are generally not concordant with definitions in Solano et al. (2022) and are ambiguously defined. Set $S$ is tied to its definition in Solano et al. (2022) by the reported size ($N=298,165$).
@@
Although Bruehl and Villarroel (2025) cited Solano et al. (2022) as the source of its primary dataset, $R$ was not used but instead the set $V$ from Villarroel et al. (2025a), based on the reported size ($N=107,875$).
Our study makes use of the following SPF datasets, all of which are publicly available: (i) set $W$ ($N=171,753$), consisting of SPFs near NeoWISE object locations [^3] http://svocats.cab.inta-csic.es/vanish-neowise/; (ii) set $R$, the most highly vetted subset from Solano et al. (2022) [^4] http://svocats.cab.inta-csic.es/vanish-possi/; and (iii) set $M$, which contains the nearly 90 million SPFs recorded in the POSS1-E and POSS1-O plates by the MAPS project (Pennington et al., 1993; Cabanela et al., 2003) [^5] https://aps.umn.edu/catalog/download/. MAPS discarded all features that did not appear in both the E (red band) and O (blue band) plates, in order to avoid recording scratches and other plate artifacts; the Galactic plane was also occluded ($|b|>20^{\circ}$) to avoid features that blend together in crowded fields. We chose the MAPS project digitization for comparison because it was independent of the DSS and SuperCOSMOS digitizations and because it records only very confident detections of celestial objects. We have also used the POSS1-E plate metadata (STScI, 2026) [^6] https://gsss.stsci.edu/skysurveys/SurveyPlateList.txt, as well as POSS1-E images [^7] https://archive.stsci.edu/cgi-bin/dss_form in DSS [^8] https://archive.stsci.edu/dss/acknowledging.html and SuperCosmos scans [^9] http://www-wfau.roe.ac.uk/sss/ (shown in Figures 2 and 9) and Pan-STARRS images [^10] https://catalogs.mast.stsci.edu/panstarrs/ (shown in Figure 2). Finally, we use the three sources cited in Bruehl and Villarroel (2025) for nuclear test dates (NNSS, 2023; Wikipedia, 2025; CHERCA, 2025).
Figure 1 illustrates the relationship between the $A$, $S$, $W$, $P$, $V$, and $R$ datasets defined in Table 1, which are defined or implied in Solano et al. (2022), Villarroel et al. (2025a), and Bruehl and Villarroel (2025). This also illustrates the ambiguity (dashed line and “?”) concerning the amount of overlap between the publicly available set $R$, which is the most vetted Solano et al. (2022) dataset, and $V$, which has not been released as of the time of this writing. None of the above papers describe the relationship between $R$ and $V$. We stress that access to $V$ is not necessary for our evaluation of the analysis and claims in Villarroel et al. (2025a) and Bruehl and Villarroel (2025).
III.2 Principal uses and limitations